a chemist has two solutions of h2so4 one has a 40% concentration and the other has a 25% concentration how many liters of each solution must be mixed to obtain 62 liters of a 31% solution
Answer by wilft1(217)

separate the information
2 solutions, we'll call them x and y
x + y = 62 liters
.4x + .25y = .31% solution
multiply 62 liters by .31 percent solution
.4x + .25y = 19.22
lets get rid of the decimal points by moving them over...
x + y = 62
40x + 25y = 1922

now, lets do this using the subtraction method, multiply the top line by - 25
then subtract to the bottom equation
-25x - 26y = -1550
40x + 26y = 1922
15x = 372
divide both sides (of the bottom equation) by 15
x = 24.8 liters
now that we know what x is, we plug that into the original equation and find out what y is...
24.8 + y = 62
y = 37.2 liters
solution 40% solution will be 24.8 liters, 25% solution will be 37.2 liters
hope this helps :)